OpenStudy (anonymous):

Subtract and write the answer in simplest form. 2 and 9/10 - 2 and 7/10

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

I'll get you started: \(2\frac{9}{10}-2\frac{7}{10}\) We would first convert to \(\frac{2*10+9}{10}-\frac{2*10+7}{10}\) \(=\frac{29}{10}-\frac{27}{10}\) =?